OpenStudy (anonymous):

Find the area inside the polar circle r=-4sin(t) and outside the circle r=2.

OpenStudy (anonymous):

I know I have to use the following equation, but I'm not really sure how to find a and b. \[\int\limits_{a}^{b}\frac{ 1 }{ 2 }(f(t)^2 - g(t)^2)\]

OpenStudy (anonymous):

\[\int\limits_{a}^{b}\frac{ 1 }{ 2 } ((-4\sin \theta)^2-2^2)d \theta\] =\[\int\limits_{a}^{b}\frac{ 1 }{ 2 } ((16\sin ^2 \theta)-4)d \theta\]

OpenStudy (anonymous):

=\[ -2 \theta + \int\limits_{a}^{b}\frac{ 1 }{ 2 } (16\sin^2 \theta)d \theta\]

OpenStudy (anonymous):

=\[ -2 \theta + \int\limits_{a}^{b}\frac{ 1 }{ 2 } (16 \frac{(1-cos2 \theta)}{2})d \theta\]

OpenStudy (anonymous):

Well, a and b are what it takes to go around the circle once, amiright? http://www.wolframalpha.com/input/?i=polar+graph+r%3D-4sin%28t%29 So, I suppose two possible a and b values are 0 and 2pi.

OpenStudy (sirm3d):

equate the two circles to get the points of intersection \[-4\sin t = 2\]

OpenStudy (sirm3d):

\[\sin t = -1/2\]\[t=7\pi/6,11\pi/6\]

OpenStudy (anonymous):

Oh, whoops, I misinterpreted the problem. Ignore me. http://www.wolframalpha.com/input/?i=plot+polar+r%3D-4sin%28t%29%2C+r%3D2

OpenStudy (anonymous):

Hm. I'll give that a shot - thank you.

OpenStudy (sirm3d):

\[A=\int\limits_{7\pi/6}^{11\pi/6}\frac{ 1 }{ 2 }\left[ \left( -4 \sin t \right)^2-\left( 2 \right)^2\right] dt\]

OpenStudy (anonymous):

So in the long run I keep coming up with the following as my solved integral, but I think I'm doing something wrong somewhere. I"m really excellent at making basic algebra errors, so I think I'm probably dropping a 1/2 somewhere... \[2\theta - 2\sin2\theta \]

OpenStudy (sirm3d):

\[16\sin^2 \theta - 4=8(2\sin^2 \theta)-4=8(1- \cos 2\theta)-4=4-4\cos 2\theta\]

OpenStudy (anonymous):

I figured it out! Thanks so much for all your patience.

OpenStudy (sirm3d):

yw.